Suppose P = 55 – 1/100 Qt and we have three identical firms. Each firm has a cost structure of $11/unit with $5,000 in fixed costs.

a. \Find the Cournot solution (Quantities, price and profits)
Answer #1

For n firms

quantity by each firm = (a - m)/b(1 + n) and price = (a + nm)/(1 + n)

Where m = marginal cost and a = vertical intercept of demand. Also b is the slope of demand function

Here a = 55 , b = 1/100 or 0.01, m = 11 and n = 3

Each firm produces q = (55 - 11)/(0.01*4) = 1100 units

Market price = (55 + 3*11)/(4) = $22

Profit = revenue - cost = 22*1100 - 1100*11 - 5000 = $7100